OpenStudy (anonymous):

Find the product of (2x + 5y)(2x − 5y)

OpenStudy (anonymous):

(a+b)(a-b)=a^2-b^2 ans=4x^2-25y^2

OpenStudy (anonymous):

4x^2-25y^2

OpenStudy (anonymous):

thank you very much